OpenStudy (anonymous):

Please help.. In the savannah, a sinkhole opened up catching some ostrich and cheetah. There was a total of 13 animals to be rescued. If 44 individual feet were pulled from the muck, how many ostrich and how many cheetahs were rescued? Set up the problem, explain why you set it up that way, and provide the solution.

OpenStudy (anonymous):

Cheetahs have 4 feets, Ostriches has 2 feets. let x = number of cheetahs, y = number of ostriches 4x + 2y = 44 x + y = 13 lets solve for number of cheetahs, they're cooler. y = 13 - x sub into original equation... 4(x) + 2(13 - x) = 44 4x + 26 - 2x = 44 2x + 26 = 44 2x = 18 x = 9 There are 9 cheetas and all of the ostriches are dead. but really. since x + y = 13 and x is 9, 13 - 9 = y 9 cheetahs, 4 ostriches

OpenStudy (anonymous):

@tonkatwuck Thank you so much!!